OpenStudy (anonymous):

Will Medal and Fan! Please Help 45 (divided by) (-1/5) A. –9 B. –225 C. 225 D. 9

OpenStudy (xapproachesinfinity):

just do what they asked you \[\frac{45}{\frac{-1 ~~}{5}}\]

OpenStudy (anonymous):

i dont understand how to

OpenStudy (anonymous):

Wait I just tried it and got -9

OpenStudy (xapproachesinfinity):

hmm what do you do if you have something like \[\frac{a}{\frac{b}{c}}\]

OpenStudy (anonymous):

divide b and c by a?

OpenStudy (xapproachesinfinity):

hmm how did you get -9?

OpenStudy (anonymous):

I divided.

OpenStudy (xapproachesinfinity):

do you know that dividing by a number is the same as multiplying with the reciprocal?

OpenStudy (xapproachesinfinity):

that was wrong

OpenStudy (xapproachesinfinity):

what is the reciprocal of -1/5?

OpenStudy (anonymous):

-5/1 I think?

OpenStudy (xapproachesinfinity):

yes! no multiply that by 45

OpenStudy (xapproachesinfinity):

now*

OpenStudy (anonymous):

i got -225

OpenStudy (xapproachesinfinity):

sounds good

OpenStudy (xapproachesinfinity):

whenever you have \[\frac{a}{\frac{b}{c}}\] you can rewrite is like \[a\times \frac{c}{b}\] those are equivalent

OpenStudy (anonymous):

Thank you so much! I appreciate that a lot :)

OpenStudy (xapproachesinfinity):

no problem